# Multipoint distributions of the KPZ fixed point with compactly supported initial conditions

### v3

@@ 2.2 An equivalent series expansion formula
Due to the block diagonal structure of the kernel $\mathbf{K}_{1}$ and $\mathbf{K}_{\mathfrak{h}}$, the Fredholm determinant $\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$ admits a series expansion, which we will be working with more frequently in the subsequent sections. To introduce the formula, we first introduce a few notations. Given $W=(w_{1},\ldots,w_{n})\in\mathbb{C}^{n}$ and $W^{\prime}=(w_{1}^{\prime},\ldots,w_{m}^{\prime})\in\mathbb{C}^{m}$, we denote
@@
which is the usual Cauchy determinant. The Cauchy determinant $\mathrm{C}(W\sqcup W^{\prime};\widehat{W}\sqcup\widehat{W}^{\prime})$ is defined in the same way with the combined variables $W\sqcup W^{\prime}$ and another set of variables $\widehat{W}\sqcup\widehat{W}^{\prime}$ with the same dimension.
@@ 3.1 Characteristic function of the initial condition
The following characterization of the initial condition comes from [Liu22], see Proposition 2.13 and the subsequent discussion in that paper for further details.
@@
$\mathbb{P}(G_{k+1}=x|G_{k}=y):=\frac{1}{2^{y-x}}\mathbf{1}_{x<y},$ (3.3)
and $\tau$ is the hitting time of $G$ to the strict epigraph of $Y$, namely
@@
$\mathrm{ch}_{Y}(v,u)$ $=\sum_{z\geq y_{N}+N}(2u+2)^{z}\cdot\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\cdot\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right].$ (3.6)
Now note that
$\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\cdot\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right]=\sum_{k=0}^{N-1}\sum_{w=y_{k+1}+1}^{y_{k}-1}\frac{2}{(2v+2)^{w+1}}\cdot\left(\frac{-v}{v+1}\right)^{k}\mathbb{P}_{G_{0}=z}[\tau=k,G_{\tau}=w],$
which is clearly analytic in $v$ for $v\in\Omega_{\mathrm{R}}$ as a finite sum of functions analytic in $v$. Moreover, since
$y_{N}+1\leq G_{\tau}\leq z,\qquad\left|\frac{-v}{v+1}\right|<1,\qquad|2v+2|>1,$
for any $0\leq\tau\leq N-1$ and $v\in\Omega_{\mathrm{R}}$, we have
$\begin{split}\left|\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right]\right|&\leq\frac{2}{|2v+2|^{z+1}}\sum_{k=0}^{N-1}\sum_{w=y_{k+1}+1}^{y_{k}-1}\mathbb{P}_{G_{0}=z}[\tau=k,G_{\tau}=w]\\ &\leq\frac{1}{|2v+2|^{z}}.\end{split}$ (3.7)
Thus by the estimates $|2(u+1)|<1<|2(v+1)|$ for $u\in\Omega_{\mathrm{L}}$ and $v\in\Omega_{\mathrm{R}}$ we have
$\left|\sum_{z\geq y_{N}+N}(2u+2)^{z}\cdot\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\cdot\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right]\right|\leq\sum_{z\geq y_{N}+N}\frac{|2u+2|^{z}}{|2v+2|^{z}},$ (3.8)
@@
$2^{\hat{z}}\cdot\mathbb{E}_{\hat{G}_{0}=z}\left[-{\hat{G}_{\hat{\tau}}-\hat{y}_{i-1}-1\choose i-\hat{\tau}-2}\frac{1}{2^{\hat{G}_{\hat{\tau}}}}\mathbf{1}_{\hat{\tau}<i-1}\right]=-\mathbf{1}_{\hat{z}\geq\hat{y}_{i-1}+{i-1}}{\hat{z}-\hat{y}_{i-1}-1\choose i-2}$,
@@ 4.3 Convergence of the series expansion
$|\log(1-z)+\sum_{k=1}^{n-1}\frac{z^{k}}{k}|\leq\frac{|z|^{n}}{n(1-|z|)},\quad\forall|z|<1,$
@@ 5.1 Density and Approximation
$\lim_{n\to\infty}\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}^{n}}(z_{1},\ldots,z_{m-1})=\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1}).$ (5.2)
@@
$\mathfrak{h}^{n}\to\mathfrak{h}\ \text{in }\mathrm{UC},\quad\text{as }n\to\infty,$
@@
where $\omega_{1}>\cdots>\omega_{k}$. To lighten the notations, we will fix $\eta,\xi$ and denote
@@
Thus $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}^{\omega_{1}}$ for all $\omega\leq\omega_{1}$. We proceed by induction. Assume $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}^{\omega_{1}}$ holds for all $\omega\geq\omega_{i}$. Now let $\omega\in[\omega_{i+1},\omega_{i})$. We will show $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}^{\omega_{i}}$, which is equal to $\chi_{\mathfrak{h}}^{\omega_{1}}$ by the induction hypothesis. By the same argument as above we have
@@
For general $\mathfrak{h}\in\mathrm{UC}_{c}$ supported on [L,L][-L,L], bounded above by $\beta$, we use Proposition (5.2) to find a sequence $\{\mathfrak{h}^{n}\}_{\geq 1}\subset m\mathrm{NW}$ bounded above by $\beta$ whose supports are contained in [L,L][-L,L] , such that $\mathfrak{h}^{n}\to\mathfrak{h}$ in $\mathrm{UC}$. Then by Proposition 5.1 we know $\chi_{\mathfrak{h}^{n}}\to\chi_{\mathfrak{h}}$ as $n\to\infty$. A minor variant of Proposition 5.1 with the same proof shows that $\chi^{\omega}_{\mathfrak{h}^{n}}\to\chi^{\omega}_{\mathfrak{h}}$ as $n\to\infty$ for any $\omega\in\mathbb{R}$ as well. Thus,
@@
We apply Proposition 4.2 to the right-hand side of (5.8) and compare the resulting formula with the right-hand side of (1.6). In order to prove Theorem 1.2 holds for $\mathfrak{h}\in m\mathrm{NW}$, it suffices to show that
@@
Note that $\mathbf{B}(\mathsf{x})\leq\mathfrak{h}(\mathsf{x}+\alpha)+\beta\iff\widehat{\mathbf{B}}(\mathsf{x}+\alpha)\leq\mathfrak{h}(\mathsf{x}+\alpha)$ where $\widehat{\mathbf{B}}(\mathsf{x}):=\mathbf{B}(\mathsf{x}-\alpha)-\beta$. The invariance of the Brownian motion implies that
@@ 6 Reduction in the equal-time case
The goal of this section is to simplify the multipoint formula (1.6) under the additional assumption that all the time parameters $\tau_{1},\ldots,\tau_{m}$ are the same, say, equal to $1$. We will show in Section 6.1 that one can get rid of the additional contour integrals with respect to the parameters $z_{i}$’s in (1.6), and get a Fredholm determinant formula for the equal-time multipoint distribution. Then in Section 6.2, we will show that the Fredholm determinant formula is equivalent to the path integral formula in [MQR21].
@@ 6.1 A new Fredholm determinant formula for the equal-time multipoint distribution of the KPZ fixed point
The following Fredholm determinant formula holds for the KPZ fixed point with a general initial condition $\mathfrak{h}$ that is compactly supported. We remark that an analogous version for the narrow wedge initial condition was obtained in [LO25, Proposition 2.1], with a slight modification on the operator space.

### v4

@@ 2.2 An equivalent series expansion formula
Due to the block diagonal structure of the kernel $\mathbf{K}_{1}$ and $\mathbf{K}_{\mathfrak{h}}$, the Fredholm determinant $\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$ admits a series expansion, which we will be working with more frequently in the subsequent sections. To introduce the formula, we first introduce a few notation. Given $W=(w_{1},\ldots,w_{n})\in\mathbb{C}^{n}$ and $W^{\prime}=(w_{1}^{\prime},\ldots,w_{m}^{\prime})\in\mathbb{C}^{m}$, we denote
@@
which is the usual Cauchy determinant. The Cauchy determinant $\mathrm{C}(W\sqcup W^{\prime};\widehat{W}\sqcup\widehat{W}^{\prime})$ is defined in the same way with the combined variables $W\sqcup W^{\prime}$ and another set of variables $\widehat{W}\sqcup\widehat{W}^{\prime}$ with the same dimension as $W\sqcup W^{\prime}$.
@@ 3.1 Characteristic function of the initial condition
The following characterization of the initial condition comes from [Liu22], see Proposition 2.13 and the subsequent discussion in that paper for further details. It claimed that the initial condition is encoded in any function satisfying two conditions in the multipoint distribution formula of TASEP. We summarize these two conditions and introduce the concept of characteristic function below.
@@
$\mathbb{P}(G_{k+1}=x\mid G_{k}=y):=\frac{1}{2^{y-x}}\mathbf{1}_{x<y},$ (3.3)
and $\tau$ be the hitting time of $G$ to the strict epigraph of $Y$, namely
@@
We remark that for the narrow wedge initial condition $y_{i}=-i$, $1\leq i\leq N$, the characteristic function defined above is equal to $\mathrm{ch}_{Y}(v,u)=\sum_{z\geq 0}(2u+2)^{z}\cdot\frac{2}{(2v+2)^{z+1}}=\frac{1}{v-u}$, which matches the corresponding characteristic function in [Liu22] for the step initial condition.
@@
$\begin{split}\mathrm{ch}_{Y}(v,u)&=\sum_{z\geq y_{1}+1}^{\infty}(2u+2)^{z}\frac{2}{(2v+2)^{z+1}}+\sum_{z=y_{N}+1}^{y_{1}}(2u+2)^{z}\cdot\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\cdot\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right]\\ &=\left(\frac{u+1}{v+1}\right)^{y_{1}+1}\cdot\frac{1}{v-u}+\sum_{z=y_{N}+1}^{y_{1}}(2u+2)^{z}\cdot\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\cdot\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right],\end{split}$ (3.6)
@@
$2^{\hat{z}}\cdot\mathbb{E}_{\hat{G}_{0}=\hat{z}}\left[-{\hat{G}_{\hat{\tau}}-\hat{y}_{i-1}-1\choose i-\hat{\tau}-2}\frac{1}{2^{\hat{G}_{\hat{\tau}}}}\mathbf{1}_{\hat{\tau}<i-1}\right]=-\mathbf{1}_{\hat{z}\geq\hat{y}_{i-1}+{i-1}}{\hat{z}-\hat{y}_{i-1}-1\choose i-2}$,
@@ 4.3 Convergence of the series expansion
$\left|\log(1-z)+\sum_{k=1}^{n-1}\frac{z^{k}}{k}\right|\leq\frac{|z|^{n}}{n(1-|z|)},\quad\forall|z|<1,$
@@ 5.1 Density and Approximation
$\lim_{n\to\infty}\prod_{\ell=1}^{m-1}\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}^{n}}(z_{1},\ldots,z_{m-1})=\prod_{\ell=1}^{m-1}\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1}).$ (5.2)
@@
$\mathfrak{h}^{n}\to\mathfrak{h}\ \text{in }\mathrm{UC},\quad\text{as }n\to\infty.$
@@
where $\omega_{1}>\cdots>\omega_{k}$. To lighten the notation, we will fix $\eta,\xi$ and denote
@@
Thus $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}^{\omega_{1}}$ for all $\omega\geq\omega_{1}$. We proceed by induction. Assume for some $1\leq i\leq k$, $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}^{\omega_{1}}$ holds for all $\omega\geq\omega_{i}$. Now let $\omega\in[\omega_{i+1},\omega_{i})$. We will show $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}^{\omega_{i}}$, which is equal to $\chi_{\mathfrak{h}}^{\omega_{1}}$ by the induction hypothesis. By the same argument as above we have
@@
For general $\mathfrak{h}\in\mathrm{UC}_{c}$ supported on [L,L][-L,L], bounded above by $\beta$, we use Proposition 5.2 to find a sequence $\{\mathfrak{h}^{n}\}_{\geq 1}\subset m\mathrm{NW}$ bounded above by $\beta$ whose supports are contained in [L,L][-L,L] , such that $\mathfrak{h}^{n}\to\mathfrak{h}$ in $\mathrm{UC}$. Then by Proposition 5.1 we know $\chi_{\mathfrak{h}^{n}}\to\chi_{\mathfrak{h}}$ as $n\to\infty$. A minor variant of Proposition 5.1 with the same proof shows that $\chi^{\omega}_{\mathfrak{h}^{n}}\to\chi^{\omega}_{\mathfrak{h}}$ as $n\to\infty$ for any $\omega\in\mathbb{R}$ as well. Thus,
@@
We apply Proposition 4.2 to the right-hand side of (5.8) and compare the resulting formula with the right-hand side of (1.6). In order to prove that Theorem 1.2 holds for $\mathfrak{h}\in m\mathrm{NW}$, it suffices to show that
@@
Note that $\mathbf{B}(\mathsf{x})\leq\mathfrak{h}(\mathsf{x}+\alpha)+\beta$ if and only if $\widehat{\mathbf{B}}(\mathsf{x}+\alpha)\leq\mathfrak{h}(\mathsf{x}+\alpha)$ where $\widehat{\mathbf{B}}(\mathsf{x}):=\mathbf{B}(\mathsf{x}-\alpha)-\beta$. The invariance of the Brownian motion implies that
@@ 6 Reduction in the equal-time case
The goal of this section is to simplify the multipoint formula (1.6) under the additional assumption that all the time parameters $\tau_{1},\ldots,\tau_{m}$ are the same, say, equal to $1$. We will show in Section 6.1 that one can get rid of the additional contour integrals with respect to the parameters $z_{i}$’s in (1.6), and get a Fredholm determinant formula for the equal-time multipoint distribution. Then, in Section 6.2 we will show that the Fredholm determinant formula matches the path integral formula in [MQR21].
@@ 6.1 A new Fredholm determinant formula for the equal-time multipoint distribution of the KPZ fixed point
The following Fredholm determinant formula holds for t he KPZ fixed point with a general initial condition $\mathfrak{h}$ that is compactly supported. We remark that an analogous Fredholm determinant for the narrow wedge initial condition was obtained in [LO25, Proposition 2.1], with the kernel conjugated to $\mathbf{T}_{\mathfrak{h}}$ below and defined on a slightly different space.